OpenStudy (anonymous):

Solve the equation 7^(1+3x) = 9^(x+2) exactly in terms of natural logarithms and then use your calculator to approximate your answer to three decimals. A) 0.523 B) 0.573 C) 0.623 D) 0.673 E) 0.723

OpenStudy (anonymous):

(1+3x)ln7=(x+2)ln9 1+3x=(x+2) *3ln3/ln7 solve it for x

OpenStudy (anonymous):

Im sorry, how do I solve this?